### v1

# ASTREA: Introducing Agentic Intelligence for Orbital Thermal Autonomy

Abstract
This paper presents ASTREA, the first agentic system deployed on flight-heritage hardware (TRL 9) for autonomous spacecraft operations. Using thermal control as a representative use case, we integrate a resource-constrained Large Language Model (LLM) agent with a reinforcement learning controller in an asynchronous architecture tailored for space-qualified platforms. Ground experiments show that LLM-guided supervision improves thermal stability and reduces violations, confirming the feasibility of combining semantic reasoning with adaptive control under hardware constraints. However, on-orbit validation aboard the International Space Station (ISS) reveals performance degradation caused by inference latency mismatched with the rapid thermal cycles characteristic of Low Earth Orbit (LEO) satellites. These results highlight both the opportunities and current limitations of agentic LLM-based systems in real flight environments, providing practical design guidelines for future space autonomy.
Index Terms—  Agentic Systems, Autonomous Operations, Space Edge Computing.
1 INTRODUCTION
Recent advances in robotics and artificial intelligence have accelerated the possibilities for autonomy, particularly in space exploration missions. However, AI could potentially transform the way lunar missions and Earth-orbiting missions are operated and monitored, especially in scenarios requiring real-time decision-making as an integral part of mission operations. This technological evolution represents a paradigmatic shift toward space systems capable of operating with minimal human intervention, particularly critical in environments where communication delays preclude direct ground supervision. Numerous autonomous control initiatives have been documented, ranging from assistance in traditional subsystems to autonomous capture and landing mission planning. These developments have demonstrated the technical feasibility of delegating critical decisions to automated systems, establishing the foundation for the next generation of fully autonomous space missions. Recently, the emergence of Large Language Models (LLMs) and their integration into the new agentic paradigm have added a new dimension to autonomous decision-making capabilities, enabling the interpretation of nuances, ambiguous contexts, and non-deterministic decision-making based on objectives or historical data, approaching human-like behavior in decision processes. This semantic reasoning capability opens previously unimaginable possibilities for space autonomy, where contextual interpretation and adaptability are fundamental to mission success. However, these large models also require substantial hardware resources in their most powerful versions, which conflicts with embedded environments that must maintain radiation tolerance and operate under severe constraints regarding power consumption, size, and temperature. This dichotomy between desired capabilities and the physical limitations of space hardware represents one of the primary challenges for the practical implementation of agentic systems in flight applications. Thermal control constitutes a critical subsystem that perfectly exemplifies these challenges, as it must maintain the operational integrity of all electronic components while managing limited computational resources in real-time. The inherent complexity of balancing variable thermal loads with dynamic processing demands requires a level of decision-making sophistication that has traditionally relied on pre-programmed rules and ground-based supervision. This work addresses this gap by implementing an agentic system on flight-qualified hardware (TRL 9) currently operating in the IMAGIN-e[1] mission, where an LLM agent provides contextual recommendations for thermal control parameters while maintaining operational independence for safety-critical applications. Beyond ground validation, the system was also deployed in on-orbit experiments aboard the International Space Station (ISS), representing, to the best of our knowledge, the first demonstration of an agentic LLM-based supervisory system executed in a real flight environment.
This system, named ASTREA (Agentic System for Thermal Regulation and Embedded Adaptation), demonstrates how advanced reasoning capabilities can be integrated into real space systems, overcoming hardware limitations through a hybrid design that combines the efficiency of reinforcement learning with the interpretability of linguistic models. The experimental validation of this approach not only confirms the technical feasibility of agentic systems in real space environments but also establishes a precedent for future implementations where intelligent autonomy is fundamental to mission success. The obtained results suggest that careful integration of advanced AI technologies into flight hardware can provide sophisticated decision-making capabilities without compromising the reliability required in critical space applications.
2 State of the art in Intelligence Autonomy based on LLMs
2.1 LLMs in Space
Space Llama [2] marks a significant milestone as the first open-source LLM (Llama 3.2) deployed and operated aboard the International Space Station (ISS). Developed by Booz Allen and Meta, Space Llama is engineered to assist astronauts in scientific and technical tasks, including predictive maintenance, autonomous access to documentation, and the replacement of physical manuals.
However, its design lacks agentic behavior, as it was not intended for autonomous mission control but rather for onboard human use.
2.2 Agentic LLM systems for Space Exploration
Recent efforts to integrate Large Language Models (LLMs) into spacecraft autonomy have primarily focused on ground-based experimentation with agentic architectures. In LLMSat [3], an LLM is proposed as a high-level control system for spacecraft, functioning as a goal-driven autonomous agent overseeing general mission operations. The study explores the LLM’s role as “Pilot,” responsible for strategic decision-making and direct vehicle control. Evaluated in a simulated software environment, the system exhibited mixed cognitive performance, with a marked decline as mission complexity increased.
Similarly, AI Space Cortex [4] introduces a centralized and explainable autonomous framework designed for future planetary missions in extreme environments, such as icy moons. This system leverages LLMs for semantic reasoning, environmental perception, autonomous site selection, and dynamic strategy adaptation—operating under distinct exploration modes (”Conservative,” ”Scientific Curiosity,” and ”Adventurous”). Validated on terrestrial testbeds, AI Space Cortex highlights the relevance of resilient and interpretable autonomy in domains with severe latency and operational uncertainty.
2.3 Combined LLM and RL Agents in Hybrid Systems
Recent advances in autonomous systems have explored the integration of heterogeneous agents that combine the semantic reasoning capabilities of LLM-based agents with the adaptive control mechanisms offered by Reinforcement Learning (RL)-based agents. These hybrid architectures seek to leverage the complementary strengths of each approach: LLM-agents excel in high-level deliberation, natural language interaction, and context-aware decision support, while RL agents operate effectively in dynamic environments through policy optimization and reward-based learning. In space applications, such architectures offer promising pathways for constructing resilient and interpretable autonomous systems capable of responding to complex operational scenarios, particularly where communication delays and environmental uncertainty impose constraints on centralized supervision.
Schoepp S. et al.[5] identify three ways of integrating LLMs in RL agents, distinguishing three primary roles: agent, where the LLM directly acts as the policy making decisions; planner, where the LLM decomposes complex tasks into subtasks or intermediate goals; and reward model, where the LLM helps generate or evaluate reward signals. Following this taxonomy, Navarro T. et al. [6] propose the use of an LLM as a reward generator within a reinforcement learning framework designed to assist in planetary landing tasks, demonstrating promising results in the practical application of LLMs to complex, high-stakes decision-making environments.
However, Chen Z. et al.[7] explicitly address the problem of hallucinations generated by LLMs. They note that LLM-dominated frameworks—where the language model plays a central role in decision-making—are vulnerable to semantic errors that may compromise the safety of RL agents, even when using state-of-the-art LLMs. To mitigate this risk, they advocate for maintaining a relative independence between the LLM and the RL agent.
Building on this last proposal, the present study examines the integration of a reinforcement learning (RL) agent for onboard satellite thermal control with a language model (LLM) agent. In this system, the LLM provides suggestions for adjusting the $\alpha$ parameter of the RL model (Soft Actor-Critic) which governs its exploratory behavior. The RL agent incorporates these recommendations while maintaining operational independence from the LLM.
The constrained hardware environment of space applications necessitates the use of smaller LLM models, usually conceived for edge computing, and also introduces significant inference latency, providing additional justification for excluding the LLM from the real-time control loop. Instead, the LLM operates as a higher-level supervisory component that analyzes extended time windows and provides medium-term strategic recommendations, while the RL agent handles immediate thermal control decisions. This hybrid architecture forms the foundation of ASTREA, a system designed to combine the semantic reasoning of LLMs with the adaptive control of RL agents in constrained onboard environments. ASTREA operates asynchronously, enabling strategic supervision without compromising real-time responsiveness.
3 Experimental Design and Implementation
3.1 General design of the experiment
The experimental setup described in this study corresponds to ASTREA, a hybrid agentic system composed of two core modules: a reinforcement learning agent responsible for thermal regulation, and a language model agent that provides semantic recommendations.
Figure 1 illustrates the overall experimental design. The RL-Agent continuously monitors the board temperature and promptly adjusts resource availability on the edge node. At the end of each episode, it transmits a summary to the LLM-Agent via an asynchronous queue. This summary includes the total number of iterations, the count of iterations where the peak temperature recorded by any sensor is within one degree of the critical threshold, and the average thermal gradient observed throughout the episode.
Meanwhile, the LLM-Agent periodically checks the episode summary queue and stores the data locally. Once a predefined time window elapses, all stored summaries are processed and fed into an LLM to generate entropy coefficient ($\alpha$) value recommendations for the SAC model used by the RL-Agent.
These recommendations are transmitted back to the RL-Agent through a separate queue. During each iteration, the RL-Agent checks for updated $\alpha$ values, and if a new recommendation is detected, it is immediately applied to the model.
Fig. 1: Schematic diagram of the experiment. The RL Agent send the iteration count, how many of that iterations have been near the threshold ($1~^{\circ}\mathrm{C}$) and the average thermal gradient of the last episode. The LLM Agent answers asynchronously the new alpha suggestion
This asynchronous communication architecture enables the coexistence of two complementary agentic modalities: a real-time RL-Agent focused on immediate thermal regulation, and a higher-latency LLM-Agent capable of deeper contextual analysis. By decoupling their operational timelines, the system leverages rapid responsiveness alongside strategic modulation, enhancing both short-term control and long-term adaptability.
3.2 LLM Agent
There is only one LLM-Agent present in the system and its only task is to decide which is the best entropy coefficient ($\alpha$) of the RL-Agent, which performs the thermal control of the system.
The LLM-Agent leverages an unrefined general-purpose LLM to address the target problem. The agent’s design capitalizes on the LLM’s broad knowledge base, guiding its reasoning through carefully crafted prompt engineering. While this approach facilitates rapid implementation, it requires the agent to abstract the problem appropriately, pre-compute values that provide direct and relevant information for decision-making, and constrain the model’s scope and operational domain to ensure the problem remains manageable.
Given the above, the LLM-Agent employs a static system prompt (see Fig 2) to configure the behavior of the language model, and a dynamic user prompt (see Fig 3) to request recommendations for updating the $\alpha$ coefficient. The generation of recommendations relies on the most recent time window of episode summaries, that includes average duration, average gradient and average of percentage steps in danger zone, along with the current $\alpha$ value.
The reason for sending this information to the LLM is to provide information on the performance of the RL-agent over the time window so that the LLM-agent can apply the rules (defined in the system prompt) and, using recommended tools, determine an $\alpha$ recommendation.
Initial preliminary experiments revealed LLM response times of approximately 10 minutes in certain instances. Considering these durations, alongside the time required for the RL model to adapt to new alpha values and cool-down periods between episodes, a 60-minute time window was established for collecting episode summaries before triggering alpha value recommendations.
This window size ensures sufficient data collection for meaningful LLM recommendations while maintaining reasonable system responsiveness for thermal control applications.
Sample System Prompt for a Threshold of 60C You optimize SAC entropy coefficient ($\alpha$) to maximize CPU usage and episode duration for thermal control. It’s optimal to keep temperature just below the threshold. THERMAL LIMIT: 60C (critical threshold) UTILIZATION PATTERNS: Low utilization (under-utilization): Short scenarios (<<10 steps) or low % in danger zone (<<20%) – indicates low CPU push or violations too quick. Medium utilization: Medium scenarios (10–60 steps) or medium % in danger zone (20–60%) High utilization (optimal utilization): Long scenarios (>>60 steps) or high % in danger zone (>>60%) – good hugging near limit without violations. Additional context: Avg thermal gradient informs heating rate (high >0.1>0.1 = fast change, use for decision nuance but not primary classification). $\alpha$ STRATEGY: Low utilization $\rightarrow$ INCREASE $\alpha$ (0.4–0.8): More exploration to find higher CPU strategies without violations. Medium utilization $\rightarrow$ MODERATE $\alpha$ (0.2–0.4): Balanced approach. High utilization $\rightarrow$ DECREASE $\alpha$ (0.05–0.2): Exploit proven high-utilization policies. Mixed/uncertain $\rightarrow$ Keep current or slightly adjust ($\pm$0.05). Consider the average percentage of steps in the danger zone ($\geq$59C) when assessing risk. A high percentage (without violations) indicates optimal max CPU near the limit; treat as high utilization if episodes long. Low percentage suggests under-utilization. Briefly justify $\alpha$ choice in 1 line, then call a tool. Tool use is mandatory.
Fig. 2: System prompt used by the LLM-Agent to configure the behavior of the language model
Sample User Prompt for Window #4 Window #4: 5 scenarios METRICS: Avg duration: 5.6 steps Avg gradient: 0.1553 Avg % in danger zone: 0.7% Current $\alpha$: 0.560 Optimize for maximum CPU usage and episode duration, tolerating temperatures near the threshold. Adjust $\alpha$?
Fig. 3: User prompt used by the LLM-Agent to ask the language model for $\alpha$ updates
The LLM-agent uses four tools for making the recommendations, which are:
  • Increase exploration: Increase exploration with alphas in the upper range ($\alpha\in[0.4,0.8]$)
  • Moderate exploration: Set exploration in a more moderate range ($\alpha\in[0.2,0.4]$)
  • Decrease exploration: Decrease exploration and exploit proven efficient policies with alphas in a lower range ($\alpha\in[0.05,0.2]$)
  • Keep alpha: Keep current value of alpha.
  • Reset alpha: Sets alpha to a default value ($\alpha=0.2$).
For the agent implementation, the provider-agnostic framework developed by OpenAI, known as the OpenAI Agents SDK [8], was chosen due to its simplicity, lightweight design, and suitability for the intended use case.
3.2.1 Large Language Model
The LLM used by the LLM-agent is a specific variant of Qwen2.5[9, 10], featuring 1.54 billion parameters and quantized to 4 bits. While Qwen2.5 includes multiple model sizes and quantization levels, this configuration was selected to balance computational efficiency and reasoning performance.
Although detailed benchmarks for 4-bit quantized versions of Qwen2.5 are limited, experimental results from the Qwen3 family suggest that models maintain competitive performance at 4-bit precision, with more pronounced degradation only observed at 3-bit quantization or lower [11]. In the context of the present use case, the selected Qwen2.5 variant has demonstrated reasoning capabilities comparable to higher-precision alternatives.
On-device inference is carried out using Llama.cpp [12], a lightweight and hardware-efficient inference framework that delivers state-of-the-art performance with minimal setup. The server implementation is compatible with the OpenAI API, enabling seamless integration with the agentic framework.
3.3 RL Agent
The RL-agent used in this work was originally presented in [13], where its design, training, and validation are detailed in the context of autonomous thermal control for small satellites. The agent is based on the Soft Actor-Critic (SAC) algorithm and was trained to regulate the satellite thermal system by managing processing resources while the cores are subjected to continuous computational load generated via the stress-ng utility.
In this work, the reward function has been revised to improve convergence and better balance performance with thermal safety. Specifically, a base survival reward and a thermal safety bonus were introduced, encouraging the agent to maintain safe operating margins while optimizing CPU usage.
In addition, a new observation feature called danger ratio has been added, which quantifies the proportion of thermal sensors operating within 10% of their defined threshold. This provides the agent with a real-time measure of proximity to thermal limits, enhancing its ability to anticipate and mitigate thermal risks.
In contrast to the configuration described in the previous study [13], where the agent controlled all cores, the current setup reserves core 0 for system management tasks, operating at maximum frequency. The RL-agent now manages the remaining 15 cores, dynamically adjusting both their frequency and power state. Figure 4 illustrates this distribution of the CPU cores in the experiment.
Fig. 4: Core-0 is reserved for the agentic thermal control system and the other 15 cores are under heavy load and managed by the agents.
3.4 Hardware Platform and Computational Constraints
The experimental setup was deployed on a flight-qualified System-on-Chip (SoC) featuring a 64-bit ARM architecture with 16 Cortex-A72 cores. The platform includes 16 GB of LPDDR4 memory and operates within a dynamic frequency range between 1.0 GHz and 2.0 GHz. No dedicated hardware accelerators (e.g., GPU, NPU) are available, which imposes strict limitations on real-time inference and parallel processing.
To ensure deterministic behavior and isolate agentic control from concurrent tasks, core 0 was reserved exclusively for ASTREA’s orchestration layer. The remaining 15 cores were subjected to continuous computational stress, simulating realistic onboard conditions. This architectural constraint directly influenced the design choices described in subsequent sections, including the adoption of a quantized LLM agent and the asynchronous execution model.
3.5 Experimental Environments: Ground Lab and ISS Deployment
To validate ASTREA under realistic operational conditions, the system was evaluated in two distinct environments: a semi-controlled ground laboratory and an external payload platform aboard the International Space Station (ISS). Each environment imposed unique constraints that shaped both the system’s behavior and its architectural design.
3.5.1 Ground Laboratory Setup
The ground experiments were conducted in a mini-rack configuration equipped with active cooling via fans. Ambient temperature was regulated for 10 hours during the day but remained unconditioned overnight, introducing natural thermal fluctuations. The SoC operated under continuous computational stress, with 15 cores fully loaded and core 0 reserved for agentic orchestration.
This setup provided a stable yet dynamic environment for evaluating ASTREA’s semantic supervision capabilities under constrained but predictable conditions. The laboratory configuration allowed for fine-grained control over thermal and computational variables, facilitating reproducible experimentation.
3.5.2 On-Orbit Deployment (ISS)
The second evaluation environment was an external payload mounted on the Columbus module of the ISS. This orbital configuration exposed the system to cyclical thermal transitions driven by the station’s 90-minute orbit—approximately 45 minutes of direct solar exposure followed by 45 minutes of eclipse.
The payload included a passive heat pipe for thermal mitigation, but no active cooling mechanisms. These constraints introduced challenges for real-time inference and agentic responsiveness, particularly in relation to the rapid environmental shifts inherent to low Earth orbit.
3.6 Baseline and Evaluation Metrics
3.6.1 Baseline: SAC with adaptive $\alpha$
The baseline system employs the same RL-agent architecture described in Section 3.3, but uses the default adaptive $\alpha$ scheduling mechanism from Stable Baselines3 [14]. This implementation automatically adjusts the entropy coefficient during training according to the internal heuristics of the algorithm, without external intervention or episode-level performance feedback 5.
Unlike the proposed LLM-guided approach, the baseline system does not collect thermal episode metrics, generate performance summaries, or receive contextual parameter recommendations based on thermal safety patterns.
Fig. 5: Only the RL-agent operates in the core-0 for performing thermal control. As in the agentic version, the load is applied on the other cores.
3.6.2 Evaluation Metrics
The following metrics are used to compare the baseline and proposed systems over a 24 hours evaluation period:
  • Number of thermal violations: Count of instances where any thermal sensor exceeds the temperature threshold during the evaluation period.
  • Average episode duration: Mean duration (in time steps) of the thermal control episodes, measured over the evaluation period.
  • CPU utilization efficiency: Inverse of remaining computational capacity, accounting for frequency scaling where high-frequency cores contribute fully to utilization, low-frequency cores contribute partially, and inactive cores contribute zero.
4 Results & Analysis
4.1 Ground Experiments
The performance of ASTREA was first evaluated against a baseline RL-only controller, under semi-controlled laboratory conditions (See 3.5.1). Within this environment, ASTREA achieved longer episode durations and significantly fewer thermal violations, particularly during the initial operational phase. These improvements underscore the value of semantic modulation in enhancing RL agent behavior under constrained and partially variable conditions.
Initially, both the agentic and baseline systems exhibited comparable behavior. However, following the first entropy coefficient ($\alpha$) adjustment recommended by the LLM-based agent, the agentic implementation transitioned into markedly longer episodes of activity—a 67.2% increase in average duration during the first 4 hours compared to the baseline. In contrast, the baseline system maintained a trajectory of gradual, incremental improvement. This divergence in behavior is illustrated in Figure 6, which highlights the influence of LLM-guided parameter modulation on the reinforcement learning (RL) agent’s operational dynamics.
02002004004006006008008001,0001{,}0001,2001{,}2001,4001{,}4001,6001{,}6001,8001{,}8002,0002{,}0002,2002{,}2002,4002{,}4002,6002{,}60005050100100150150IterationEpisode Length MeanAgenticBaselineAlpha updates
Fig. 6: Episode length mean for baseline and agentic executions. The dashed red lines show the iterations in which the LLM-agent gave recommendations for $\alpha$ parameter in the agentic implementation.
Table 1 presents a comparative analysis of both systems across three key metrics: average episode duration, number of thermal violations, and average CPU utilization. The agentic system not only sustained longer episodes but also achieved a 58.5% reduction in thermal violations during the early phase, and a 42.1% reduction over 24 hours, indicating significantly improved thermal compliance. These gains were achieved with minimal impact on CPU usage, which remained nearly constant—only a 1.9% increase during the first 4 hours and a 1.5% decrease over the full 24-hour period.
Longer episodes are indicative of improved thermal stability, as the agent successfully maintains system temperature below the critical threshold for extended periods. A reduced number of thermal violations further reflects enhanced adherence to thermal constraints. These metrics are inherently interdependent, given that episodes terminate upon exceeding the thermal limit—thus, longer durations typically correlate with fewer violations.
It is important to note that the RL agent is also tasked with maximizing CPU usage without triggering thermal violations. This dual objective prevents the agent from simply throttling performance to extend episode duration, ensuring a balanced optimization strategy.
During the initial 4-hour period, the agentic system demonstrated substantial improvements in both episode duration and thermal compliance, likely due to the exploratory strategies introduced by the LLM-agent. Over the full 24-hour period, while the difference in episode duration narrowed to just 5.2%, the agentic system maintained superior thermal performance, suggesting that LLM-guided modulation yields early advantages that persist in long-term stability. This early advantage may also contribute to the system’s ability to dynamically adapt to shifting operational conditions, further reinforcing the role of the LLM-agent as a flexible and context-aware modulator in unpredictable space environments.
Table 1: Experiment results on ground in early stages and after 1 day of execution.
Metric Duration Baseline (Mean $\pm$ Std. Dev) Agentic (Mean $\pm$ Std. Dev)
Avg. Episode Duration First 4h $47.17\pm 18.15$ $\mathbf{78.83\pm 11.33}$
24h $135.24\pm 32.94$ $\mathbf{142.29\pm 8.06}$
Thermal Violations First 4h $39.33\pm 9.29$ $\mathbf{16.33\pm 2.08}$
24h $88.67\pm 20.50$ $\mathbf{51.33\pm 4.04}$
Avg. CPU Usage (%) First 4h $25.81\pm 5.00$ $\mathbf{26.30\pm 2.56}$
24h $\mathbf{16.49\pm 5.57}$ $16.24\pm 4.46$
Finally, as shown in Figure 7, LLM-agent response times ranged from approximately 40 seconds to over 8 minutes, depending on prompt size and the operational status of core 0. These latencies clearly indicate that, within the present configuration, incorporating the LLM-agent into the real-time control loop of the RL-agent is not feasible. Such delays in decision-making could pose a thermal risk to the systems under supervision, particularly during high-stress operational scenarios.
It is important to note, however, that in this experiment, all LLM processing was confined to core 0, thereby precluding any parallelization strategies typically available when leveraging multiple cores. This architectural constraint significantly contributes to the observed latency.
112233445566778899101011111212131305050100100150150200200250250300300350350400400450450500500PromptsTotal Processing Time (s)
Fig. 7: Total processing time (latency) reported by the LLM for twelve individual prompt executions.
4.2 On-Orbit Experiments (ISS)
ASTREA was subsequently deployed aboard the International Space Station (ISS) to validate feasibility in a true operational environment (See 3.5.2). In this experiment, the agentic system exhibited shorter episode durations than the RL-only baseline, but also achieved a reduction in the number of thermal violations. CPU usage showed a slight increase when the LLM-agent was active (Table 2).
Table 2: Representative results from on-orbit experiments aboard the ISS.
Metric Baseline (Mean $\pm$ Std. Dev) Agentic (Mean $\pm$ Std. Dev)
Avg. Episode Duration $\mathbf{224.26\pm 163.81}$ $146,60\pm 71.04$
Thermal Violations $204\pm 201.72$ $\mathbf{184.33\pm 69.47}$
Avg. CPU Usage (%) $18.84\pm 13.01$ $\mathbf{19.72\pm 17.43}$
The mixed performance profile can be partially explained by a temporal mismatch between the LLM decision cycle and the orbital thermal dynamics, combined with the stabilizing effect of the agentic policy on thermal control. Unlike terrestrial experiments, where ambient temperature evolves gradually and predictably, the ISS undergoes rapid and extreme thermal transitions due to its 90-minute orbital period. As the station circles the Earth, it experiences approximately 45 minutes of direct sunlight followed by 45 minutes of eclipse, resulting in sharp temperature swings across its external surfaces.
In this experiment, the analysis window provided to the LLM was shortened to 15 minutes to approximate this orbital cadence. However, as illustrated in Figure 7, the LLM required up to 8 minutes to generate a recommendation in the current configuration. During such an interval, the ISS may already undergo a complete sun/eclipse transition, rendering the suggested $\alpha$ value outdated at the moment of application. Consequently, some updates introduced exploratory behavior precisely during phases that demanded conservative exploitation, while in other cases the agentic adjustments reduced the magnitude and frequency of thermal violations.
Despite the mixed outcomes observed in orbit, the successful deployment and execution of the agentic system aboard the ISS validates the feasibility of the onboard architecture under real flight conditions. This operational confirmation represents a critical milestone for future agentic system implementations in space environments.
4.3 Comparative Discussion
The discrepancy between ground and on-orbit results underscores an important limitation of agentic LLM-based systems in flight conditions: when the environment’s dominant timescale (e.g., orbital sun/eclipse transitions) is comparable to or shorter than the LLM inference latency, semantic supervision can become counterproductive. However, the observed reduction in thermal violations suggests that, even under latency constraints, agentic policies can contribute to improved stability in certain operational metrics. This finding highlights a general principle for the deployment of LLMs in safety-critical space systems: they are unsuitable for direct real-time intervention unless the latency of decision-making is significantly smaller than the dynamics of the controlled process.
Nevertheless, these mixed results are highly informative. They provide empirical evidence that complements the successful ground experiments and establish practical design guidelines.
5 Implications for Future Space Autonomy
The experimentation confirms that the integration of linguistic models for semantic supervision of other processes is feasible in flight-qualified equipment within the transformative New Space era. Agentic systems are thus positioned as a viable and, given their inherent characteristics, transformative approach to integrating control logic into traditional subsystems. The deployment of ASTREA demonstrates how agentic systems can serve as supervisory layers for traditional control architectures. By combining linguistic reasoning with adaptive control, ASTREA exemplifies a scalable approach to onboard autonomy, capable of supporting fault detection, mission planning, and subsystem coordination.
Given the analytical capabilities that extend beyond rigid rule-based systems, it is reasonable to anticipate that the integration of agentic systems aboard spacecraft will become the norm rather than the exception in the near future. While these systems may not assume complete mission control, they are well-positioned to assist other subsystems, collaborate with additional LLM agents or other agent types as proposed in this study, or simply analyze and characterize emergency levels from satellite telemetry to provide early warnings to ground control systems. Furthermore, an agentic system could autonomously plan onboard analyses of sensor data using specialized models or corroborate telemetry through alternative methods utilizing other sensors typically present on the satellite, mirroring the approach an engineer would take when confronted with anomalous telemetry readings.
This capability assumes even greater significance in remote exploration missions, providing the ability to perform contextual analysis while applying pseudo-reasoning to decision-making processes, thereby conferring a high degree of automation and autonomy to space systems.
This type of onboard analysis and automation would be impractical to achieve with other technologies. Even implementing a fraction of such analytical capability would require extensive development and validation periods under simulated extreme conditions, inevitably excluding parameters not considered during the design and implementation phases of such software. This limitation eliminates any capacity for improvisation, which is often precisely what is needed in highly dynamic and unpredictable environments.
However, incorporating the most powerful linguistic models into flight-qualified hardware remains technically unfeasible at present. This constraint necessitates maintaining short context lengths in prompts and structuring them in such a way that smaller models, compatible with current flight hardware capabilities, can properly contextualize the information received in the prompt and respond appropriately.
Another critical consideration in the deployment of such systems is the grounding problem. As Schoepp S. et al. expose, this manifests in the integration of LLM-agents with other systems that expect more structured inputs, such as reinforcement learning agents, requiring the adaptation of natural language generated by the LLM-agent into executable actions for the RL-agent [5]. One approach to address or mitigate these grounding challenges, as implemented in this study’s use case, is to employ the well-established Tool Use design pattern. This approach constrains the model to utilize tools that generate concrete actions, enabling the model itself to perform the necessary adaptation as part of its analytical process for response generation.
6 Conclusions and Future Work
This study addresses the utilization of agentic systems in real flight equipment to provide an advanced layer of pseudo-reasoning at the edge. The research evaluates the utility of small linguistic models, which despite their contextual understanding limitations, remain feasible given current hardware constraints in flight equipment, deployed at the edge for autonomous control applications.
ASTREA represents a pioneering implementation of agentic autonomy in flight-qualified hardware. By integrating a quantized LLM with a reinforcement learning agent, ASTREA establishes a framework for semantic supervision and adaptive control in space systems. Its modular design and asynchronous architecture offer a scalable blueprint for future multi-agent deployments.
To validate this approach, we experimented with the integration of an LLM-agent with a reinforcement learning agent previously employed in earlier studies for autonomous thermal control of an edge processing node through the management of available processing resources. While the models present inherent limitations, our findings demonstrate that through the application of various optimization techniques, their utilization becomes viable, yielding clear improvements in the thermal control agent’s performance, during the cold start phase, by modifying its behavior based on obtained results. This highlights the potential of linguistic models for long-term context analysis, efficiently complementing the capabilities of reinforcement learning models in decision-making processes.
The present study employed a single LLM agent with significant response time limitations. This constraint would become even more critical in multi-agent systems where bottlenecks could easily emerge, making it worthwhile to explore these limitations in future research.
Since the flight equipment utilized lacks an accelerator for optimal LLM model execution, deployment occurred on the CPU alongside other concurrent processes. While functionality has been demonstrated, the incorporation of a space-qualified accelerator would fundamentally transform system performance, potentially enabling the use of more complex models or alternative execution designs such as shorter analysis windows and more frequent evaluations.
In the analyzed use case, the LLM agent suggests entropy coefficient values for a reinforcement learning model. However, numerous other parameters within such models could also be analyzed by the LLM agent, including the possibility of adjusting the reward function to shape the behavioral objectives of the RL agent, as explored by Navarro T. et al. [6].
Moreover, the capabilities of LLMs are not limited to fine-tuning existing parameters; their potential extends to higher-level control mechanisms. Thanks to their ability to interpret broader contexts, LLM agents could monitor instrument degradation throughout a mission and adapt usage strategies accordingly. They might also detect periodic patterns in environmental conditions based on orbital position and dynamically select among reinforcement learning models with varying characteristics—such as more conservative or more aggressive behaviors. Furthermore, these models could be combined or switched in real time depending on evolving operational priorities. For instance, the system might prioritize energy efficiency over computational performance during specific mission phases, or vice versa, depending on the mission’s current objectives.
Finally, this study selected a general-purpose model that may lack sufficient thermal control domain knowledge. Repeating the experiment with a model specialized in thermal control would be highly interesting, as enhanced task-specific knowledge could enable more grounded analysis without requiring extensive prompt engineering to condition its pseudo-reasoning capabilities.
In summary, ASTREA proved effective in the lab and, in orbit, reduced thermal violations and increased CPU utilization in line with mission goals, although episode durations were shorter. These results confirm the feasibility of LLM‑based supervision in flight and provide concrete design criteria for future spaceborne autonomous systems.
Acknowledgments
I am grateful to my colleague Julian Cobos Aparicio for his careful review of the manuscript and for providing valuable editorial suggestions that enhanced its clarity. I also thank my friend and colleague David Alami for introducing me to the field of agentic AI, which inspired the conception of this paper.